OpenStudy (anonymous):

What constant acceleration is required to increase the speed of a car from 28 mi/h to 53 mi/h in 5 s? (Round your answer to two decimal places.)

OpenStudy (lgbasallote):

use the formula \[\huge V_2 = V_1 + at\] where V2 is the final speed v1 is the initial speed a is acceleration t is time so plug in! \[\huge 53 = 28 + a(5)\] \[\huge 53 - 28 = 5a\] \[\huge 35 = 5a\] \[\huge a = 7 \;m/s^2\]

OpenStudy (shane_b):

Or \[a=\frac{V_f-V_i}{\Delta t}=\frac{53mi/h-28m/h}{5s}=?\]

OpenStudy (anonymous):

ok thanks

OpenStudy (lgbasallote):

i made a slight mistake \[\huge 53 - 28 = 5a\] \[\huge 25 = 5a\] \[\huge a = 5 \; \; m/s^2\]